OpenStudy (anonymous):

Hailey’s new puppy, Sparky, weighed 4 pounds at six weeks old. If the puppy gains 2.5 pounds each month, at how many months will Sparky weigh 14 pounds? Is this situation modeled by a linear function or an exponential function?

OpenStudy (anonymous):

ok u need an explanation, or a quick answer

OpenStudy (anonymous):

either one i just need to understand so explanation

OpenStudy (anonymous):

this is modeled ny a linear function Slope-intercept form y=mx+y-intercept the y intercept is 4 for you started with four pounds your m(slope) is 2.5 per month so 2.5x4=10 10+4=14 so 4 months to increase by 10 pounds and 1.5 months to increase by 4 so 4+1.5=5.5 months the answer

OpenStudy (anonymous):

is it clear now?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

no ;) can u medal and fan me?

OpenStudy (anonymous):

np*